OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope 4/5 and y-intercept (0, -1) and graph the line.

OpenStudy (aroub):

y=4/5x-1

OpenStudy (anonymous):

I need a break down on this

OpenStudy (aroub):

okay..

OpenStudy (aroub):

y=mx+b

OpenStudy (aroub):

m=slope

OpenStudy (aroub):

which is 4/5

OpenStudy (aroub):

b=y-intercept

OpenStudy (aroub):

which is -1

OpenStudy (aroub):

so now u put them in the form y=mx+b

OpenStudy (aroub):

y=4/5x-1

OpenStudy (anonymous):

Thank you so much